OpenStudy (anonymous):

In set builder notation, how do you write the solutions of the inequality? –2 –x ≥ –1

OpenStudy (anonymous):

-2-x>=1 -x>=1+2 -x>=3 x<=-3 says "x is less than or equal to -3" in the number line you have _________I____________I______________I______________I____________I____________I -5 -4 -3 -2 -1 0 highlighted the point x=-3 and all the points to the left of -3 thus you shade in x=-3 and the number line to the left of x=-3. in set notation you would write --------> ANSWER: (-infinity, -3] the parenthesis is to denote that the point is not included or if you are using infinity (-infinity or +infinity) and the bracket is to denote that the point is included in the set (when you have an equal sign).

OpenStudy (anonymous):

thanks so much :)